# Majorana Tensor Decomposition: A unifying framework for decompositions of fermionic Hamiltonians to Linear Combination of Unitaries

### v2

@@ Acknowledgements
We are grateful to Nathan Wiebe and Nicholas Rubin for valuable discussions. I.L. and A.F.I. gratefully appreciate financial support from the Mitacs Elevate Postdoctoral Fellowship and Zapata Computing Inc.
@@ Appendix A LCU circuits
Table 4: T-gate, number of logical qubits, and number of controlled $R_{Z}$ rotations cost of quantum circuits of interest for this work. Reusable qubits corresponds to those that are left in the |$\ket{0}$ state after the directive is executed, and is also included in the logical qubit count. We have used $b_{n}\equiv\lceil\log_{2}n\rceil$, $k_{n}\equiv\lfloor\log_{2}n\rfloor$, and $l_{n}\equiv\lceil\log_{2}(n/(2^{k_{n}}))\rceil$ such that $b_{n}=k_{n}+l_{n}$. $\mu_{n}\equiv\lceil\log_{2}(n\epsilon)\rceil$ corresponds to the number of bits required for the “alt” register used when preparing a coherent superposition of $n$ coefficients with an accuracy of $\epsilon$. In general, $\epsilon$ corresponds to the target accuracy for the given directive and $\lambda$ to the 1-norm of the LCU. $N$ represents the number of spacial orbitals. Numbers in braces for $\mathtt{PREPARE}$ circuit correspond to additional resources when circuits are controlled by an additional qubit, shown by the |$\ket{ctl}$ register in the corresponding figures. We always consider all $\mathtt{SELECT}$ operations to be controlled by a |$\ket{ctl}$ qubit. All constants are defined in the caption of the corresponding figure. As shown in Ref. 62, $R_{Z}$ rotations can be optimally implemented for a target accuracy $\epsilon<0.016$ with an average T-gate cost of $9.6783.067*\log_{2}\left(1/\epsilon\right)+$ 9.678 and one additional reusable qubit. The controlled Givens$(N)$ directive corresponds to an orbital rotation consisting of $N-1$ Givens rotations that are controlled by a register where the corresponding angles have been previously loaded, as shown in Refs. 24, 20. The rotations for this directive can be implemented using the phase gradient technique [63] for a cost no larger than $7$ T-gates per rotation. Note that our notation for $N$ differs by a factor of $2$ to that in Ref. 20.
@@ Appendix B Computational details
All electronic Hamiltonians in the minimal STO-3G basis set were obtained with the PySCF [65, 66, 67] package. Results involving tensor decompositions of two-electron integrals correspond to the lowest rank sufficient to obtain a converged two-electron tensor with the 2-norm error $\sum_{ijkl}|\Delta g_{ijkl}|^{2}$ below $1\times 10^{-6}$. All LCU decompositions were done using the QuantumMAMBO package on Julia, which is available at https://github.com/iloaiza/QuantumMAMBO.jl.
@@
For the molecules in Table 2, the nuclear configurations are:
  • R(H – H) = $0.741\textrm{\AA}$ for H2
  • R(Li – H) = $1.595\textrm{\AA}$ for LiH
  • R(Be – H) = $1.326\textrm{\AA}$ with a collinear atomic arrangement for BeH2
@@
[…]
For the hydrogen chains, the R(H – H) distance was taken to be $1.4\textrm{\AA}$.

### v3

@@ Acknowledgements
We are grateful to Nathan Wiebe and Nicholas Rubin for insightful discussions, and Alexey Uvarov, Sangeeth D. Kallullathil, Shreyas Malpathak, and Joshua T. Cantin for valuable comments on the manuscript. I.L. and A.F.I. gratefully appreciate financial support from the Mitacs Elevate Postdoctoral Fellowship and Zapata Computing Inc.
@@ Appendix A LCU circuits
Table 4: T-gate, number of logical qubits, and number of controlled $R_{Z}$ rotations cost of quantum circuits of interest for this work. Reusable qubits corresponds to those that are left in the |$\ket{0}$ state after the directive is executed, and is also included in the logical qubit count. We have used $b_{n}\equiv\lceil\log_{2}n\rceil$, $k_{n}\equiv\lfloor\log_{2}n\rfloor$, and $l_{n}\equiv\lceil\log_{2}(n/(2^{k_{n}}))\rceil$ such that $b_{n}=k_{n}+l_{n}$. $\mu_{n}\equiv\lceil\log_{2}(n\epsilon)\rceil$ corresponds to the number of bits required for the “alt” register used when preparing a coherent superposition of $n$ coefficients with an accuracy of $\epsilon$. In general, $\epsilon$ corresponds to the target accuracy for the given directive and $\lambda$ to the 1-norm of the LCU. $N$ represents the number of spacial orbitals. Numbers in braces for $\mathtt{PREPARE}$ circuit correspond to additional resources when circuits are controlled by an additional qubit, shown by the |$\ket{ctl}$ register in the corresponding figures. We always consider all $\mathtt{SELECT}$ operations to be controlled by a |$\ket{ctl}$ qubit. All constants are defined in the caption of the corresponding figure. As shown in Ref. 65, $R_{Z}$ rotations can be optimally implemented for a target accuracy $\epsilon<0.016$ with an average T-gate cost of $9.6783.067*\log_{2}\left(1/\epsilon\right)+$ 9.678 and one additional reusable qubit. The controlled Givens$(N)$ directive corresponds to an orbital rotation consisting of $N-1$ Givens rotations that are controlled by a register where the corresponding angles have been previously loaded, as shown in Refs. 24, 20. The rotations for this directive can be implemented using the phase gradient technique [66] for a cost no larger than $7$ T-gates per rotation. Note that our notation for $N$ differs by a factor of $2$ to that in Ref. 20.
@@ Appendix B Computational details
All electronic Hamiltonians in the minimal STO-3G basis set were obtained with the PySCF [68, 69, 70] package. Results involving tensor decompositions of two-electron integrals correspond to the lowest rank sufficient to obtain a converged two-electron tensor with the 2-norm error $\sum_{ijkl}|\Delta g_{ijkl}|^{2}$ below $1\times 10^{-6}$. All LCU decompositions were done using the QuantumMAMBO package on Julia, which is available at https://github.com/iloaiza/QuantumMAMBO.jl. All least square optimizations are performed using the Julia package LeastSquaresOptim.jl.
@@
Details of THC implementation
The two-body tensor of the Hamiltonian $g_{ijkl}$, i.e., the terms of the Hamiltonian quartic in Majorana operators, for a system with $N$ spatial orbitals, is decomposed using
$g_{ijkl}=\sum_{\mu\nu}\Lambda_{\mu\nu}\left[\vec{U}_{\mu}\otimes\vec{U}_{\mu}^{T}\right]_{ij}\left[\vec{U}_{\nu}\otimes\vec{U}_{\nu}^{T}\right]_{kl}$ (67)
where $\Lambda$’s are related to the $\zeta$’s from Eq. (58) as $\Lambda_{\mu\nu}=4\zeta_{\mu\nu}$ and $\vec{U}$’s are $N$ dimensional normalized vectors.
To see that Eq. (67) follows from Eq. (58), one needs to rewrite Eq. (58) as
$\hat{H}=\sum_{\mu\nu}\sum_{\sigma\tau}\zeta_{\mu\nu}\left(\hat{U}_{\mu}^{\dagger}(2\hat{n}_{1\sigma}-1)\hat{U}_{\mu}\right)\left(\hat{U}_{\nu}^{\dagger}(2\hat{n}_{1\tau}-1)\hat{U}_{\nu}\right)+\hat{h}_{1e}$ (68)
where $\hat{n}_{1\sigma}$ and $\hat{n}_{1\tau}$ are occupation number operators and $\hat{h}_{1e}$ contains only one-electron and constant terms. Suppressing the spin-indices for clarity, we can rewrite the above expression as
$\hat{H}=\sum_{\mu\nu}\Lambda_{\mu\nu}\left(\hat{U}_{\mu}^{\dagger}\hat{n}_{1}\hat{U}_{\mu}\right)\left(\hat{U}_{\nu}^{\dagger}\hat{n}_{1}\hat{U}_{\nu}\right)+\tilde{h}_{1e},$ (69)
where remaining constant terms have been absorbed into $\hat{h}_{1e}$ to give $\tilde{h}_{1e}$. Since the one-electron terms can be treated optimally with DF, we focus on the first term of the RHS. We exploit the isomorphism between Lie algebra of one-electron excitation operators for a given spin and $N\times N$ matrix algebra, correspondingly, $\left(\hat{U}_{\mu}^{\dagger}\hat{n}_{1}\hat{U}_{\mu}\right)$ maps to $\mathbb{U}_{\mu}^{\dagger}\mathbb{n}_{1}\mathbb{U}_{\mu}$, where $\mathbb{U}_{\mu}$ and $\mathbb{n}_{1}$ are $N$-dimensional square matrices. The elements of $\mathbb{n}_{1}$ are given by $\left[\mathbb{n}_{1}\right]_{pq}=\delta_{pq}\delta_{p1}$. It is straightforward to show that $\left[\mathbb{U}_{\mu}^{\dagger}\mathbb{n}_{1}\mathbb{U}_{\mu}\right]_{pq}=\left[\mathbb{U}_{\mu}\right]_{1p}\left[\mathbb{U}_{\mu}\right]_{1q}$. Since each row and column of a unitary matrix is a normalized vector, we can write $\left[\mathbb{U}_{\mu}\right]_{1p}$ as the $p^{\rm th}$ component of a normalized vector $\vec{U}$. Thus, we can rewrite the first term in the RHS of Eq. (69) as the RHS of Eq. (67). To arrive at the LHS of Eq. (67), the terms in the original Hamiltonian $\hat{H}$ quartic in fermionic creation and annihilation operators are considered. We note that the fermionic excitation operator $\hat{E}_{q}^{p}=\hat{a}_{p}^{\dagger}\hat{a}_{q}$ is isomorphic to the matrix $\mathbb{E}_{q}^{p}$ whose elements are $\left[\mathbb{E}_{q}^{p}\right]_{ij}=\delta_{ip}\delta_{jq}$. Mapping $\sum_{pqrs}g_{pqrs}\hat{E}_{q}^{p}\hat{E}_{s}^{r}$ to $\sum_{pqrs}g_{pqrs}\mathbb{E}_{q}^{p}\mathbb{E}_{s}^{r}$ and then exploiting the isomorphism we arrive at Eq. (67). The norm of the THC decomposition is $\sum_{\mu\nu}\absolutevalue{\Lambda_{\mu\nu}}$ | in addition to $\Delta E/2$ of $\tilde{h}_{1e}$.
For a system with $N$ spatial orbitals, we start with $K$ $\vec{U}$’s. All $\Lambda$’s are initialized to unity and each component of each $\vec{U}$ is initialized to $1/\sqrt{N}$. We choose the target accuracy $\epsilon$ to be a residual 2-norm of $1\times 10^{-6}$. The cost function for the THC decomposition is
$\sum_{pqrs}\left(g_{pqrs}-\sum_{\mu\nu}\Lambda_{\mu\nu}U_{\mu p}U_{\mu q}U_{\nu r}U_{\nu s}\right)^{2}+\sum_{\mu\nu}\left(\alpha\absolutevalue{\Lambda_{\mu\nu}}^{\beta}\right)^{2}$ (70)
where $\alpha$ and $\beta$ are hyper-parameters that tune relative importance of the two contributions. The first term ensures the convergence of the THC ansatz to the original two-body tensor, while the second term penalizes large 1-norm. We found that the optimal decompositions are obtained via the choice of parameters $\alpha=0.2$ and $\beta=1$. The form of the cost function allows us to optimize for $\Lambda$’s and $\vec{U}$’s via non-linear least square optimization, which we carried out using the Levenberg-Marquardt algorithm [57, 58]. If the 2-norm of the residual two-body tensor is greater than the desired accuracy $\epsilon$, we decompose the residual two-body tensor using another set of $K$ normalized vectors. The initialization of the new $\Lambda$’s and $\vec{U}$’s is the same as before. This step is repeated until the desired accuracy is achieved. A total of $J$ sets each containing $K$ normalized vectors are required to achieve the desired accuracy. We worked with $K=3N$ for the H-chains reported in Fig. 5 and Table 3, and $K=4N$ for the small molecules reported in Table 2, with the exception of NH3 for which $K=36$. For all THC decompositions reported in this paper, we found that $J=2$ was sufficient to reach the target accuracy. In order to estimate the resources required to encode the THC decomposition into a quantum circuit, we have assumed that the allowed error in the quantum phase estimation is 1 mili-Hartree and the number of bits used to represent the LCU coefficients and rotation angles are 10 and 16, respectively.